A total of juniors and seniors were given a math test. The 35 juniors attained an average score of 80 while the 15 seniors obtained an average of 70 . What was the average score for all 50 students who took the test ?
Found 2 solutions by Boreal, satyareddy22:
Answer by Boreal(15235)

35*80=2800
15*70=1050
Sum is 3850
number is 50
Average is 77
average is total/n
Total is average *n
You use the second to get the total, then once the total is obtained, divide by 50 to get the overall average.

Answer by satyareddy22(84)

The 35 juniors attained an average score of 80 = 80 *35 = 2800
The 15 seniors obtained an average of 70 = 70 * 15 = 1050
The total score for all 50 students who took the test = 2800 + 1050 =3850
The average score for all 50 students who took the test = 3850/50 = 77